“breve” stato dell’apparato: risultati e problemi anton, ian, nicolas, saleh, salvatore barrel...

Download “Breve” stato dell’apparato: Risultati e Problemi Anton, Ian, Nicolas, Saleh, Salvatore Barrel Endcap Endcap Endcap Cooling Endcap Cooling Last Minute

Post on 18-Jan-2016

214 views

Category:

Documents

0 download

Embed Size (px)

TRANSCRIPT

Slide 1

Breve stato dellapparato:Risultati e ProblemiAnton, Ian, Nicolas, Saleh, Salvatore Barrel Endcap Endcap Cooling Last Minute .

Domenica !!?..neanche a bergamo alta

BarrelHV&LV problemsLV barrel: Lost of threshold control due to FEB or DB problem - near 3% of chambers per wheel (apart from W+1 which is fine)HV barrel: several chambers in SG (~3-4), several cables from the cable chain broken (~10), several P.S distributors broken (~4-5), several chamber with high current (~2-3). A dedicated excel file is being prepared with all HV problem types. Once available and updated all the numbers and exact names will be known.Noise Study:Performed ActionsHardware InterventionsStar washersMultiple groundingCopper tape (for better connectivity to ground)Additional RPC ShieldingExternal HV&LV systemImproving grounding (capacitors + shortening LVDS shielding cables + additional ground to iron)LV power supply instabilities checked (7 modules repaired)Software InterventionsReducing noise by setting threshold and maskingNoise ProblemsThere are two main problems:Coherent noiseNoise Fluctuation in time

Still to be doneTest new procedure to set thresholdsTest new masking procedure

Coherent Noise

Noise FluctuationsConclusionsProbably there is no good contact in all ground connection on the front panel since copper tape and alodine improve the chamber performanceProbably there is no good contact between the aluminum covers and the preloaded bars. M.Jonson suggested that capacitive coupling between the top covers and the FEB is the reason for introducing the noiseLV power supply instability sensitive to external disturbance. Good candidate source for coherent noise Very large spread in threshold settings that could lead to chip self-excitation. Good candidate source for noise fluctuationsQuestions/DiscussionDo we monitor the chambers with installed star washers? Do we see any difference?Do we monitor the chambers with additional AL covers? Do we see any difference?Do we think yet that the LV power supply instability is sensitive to external disturbance and could be a good candidate source for coherent noise?Do we know the effect of the shortening of LVDS cables? Can we conclude that the new procedure to set thresholds have solved our problems in noise fluctuations?Trigger InterventionsChange RJ45 connectors in order to improve grounding of CCU rings (14 tower + 4 already done)Upgrade CBChange LAN Trafos with new ones with magnetic shieldingFix known problems YEN1nearInvestigate one tower failure during CRAFT09 RB+1farActions do be doneTerminate the story of swapped signal cablesAction matrix to be fulfilledSpare list & Spare managementProcedure and names for hardware interventions during LHC

EndcapHigh Voltage Preparation of the high voltage system for all 3 negative side RPC stations.- Assembly and testing of the HV distribution boxes. (12 for negative Endcap RPCs)- Installation of hardware in USC. (4 EASY crates, 4 fan units,18 modules)- Cables connection from DBs to modules. (120 cables)- Preparation of the Jupiter connectors , connection with DBs and labelling USC.(480 connectors)- Patch Panel preparation in UXC. (grounding, routing and connection of 216 cables) HV racks temperatures sensors installation and connection with PP. (12 T sensors).

Problems appeared and present status: In Total 33 channels tripping problems were seen in both +ve & -ve Endcap RPCs. 23 have been fixed by changing the connector at PP, chambers and DBs sides. 10 still need to be fixed but have been minimized by disconnecting the gaps.-In Total 16 channels spikes observed 09 have been fixed by changing connectors at PP, chambers and DBs sides. 07 still need to be fixed but have been minimized by disconnecting the gaps. One side labels from module to EASY crates and routing of the cables in cables tray still to be done.

Low Voltage Preparation of the Low voltage system for all 3 negative side RPC stations.Installation of hardware in UXC. (6 AC/DC convertor, 6 EASY Crates , 18 LV boards)Daisy chain preparation of DCS cables from MAO, FEBs , LBBs crates and then to next rack .(6 racks are in use for Z side).Cables connection with LV boards, reparation of APP45 connectors and cables routing . (432 connectors , 108 cables).Preparation of APP75 for 48V from MAO to Easy crates.(24 plugs)Preparation of flat ribbon cables with 50 pins connectors for racks internal daisy chain. (24 cables ).Installation of branch controllers and mapping of the hardware.

LV Problems:Negative side ~9 channels tripping problems were seen due to bad connectors of APP45: all have been fixed.One MAO broken replaced with spare.In Total 8LV boards gave problems during last 5 months in whole endcap RPCs system. All repaired from EP now every thing is fine.Two new cables are pulled from chambers to racks in RE+3 and RE-1.

Readout SystemConnections at both side (chambers and link board side) already done during pre-commissioning.

For commissioning during connectivity and threshold scan test following problems observed.

- 6 faulty FEBs replaced with new ones from different stations. - I2C problems seen in some chambers, many solved but 4 still to be done. - Many swap of signal cables seen at LBBs and Chambers level all have been solved.Gas SystemGas leak search performed at all stations.In 3 sectors open gas circuits detected which have been fixed. (2 in RE-3 and 1 in RE-1)In 5 sectors big leak (500 to 1300ml/h in approx. 18 L volume) is observed . In 4 sectors leaks have been fixed. One at RE+2 still to be cured (problem in access time).Modified 6 chambers of one gas sectors from one gas in/out to two in/out at RE+2 and made this sector semi parallel 10degree.Total 6 sectors made semi parallel with respect to gas by using custom gas distribution boxes at RE+2.Bad calibrated gas flow cells have been changed and cells have been equalized.Gas racks flow cells have been labelled. 64 Gas Humidity sensors have been installed on gas pipes (48 in the barrel)87 T sensors have been installed

Temperature/Humidity sensorsCampaign to equip the Endcap RPCs with temperature sensors have been almost ended as the original plan (1 sensor each 6 chambers). T is a key element to monitor. Additional sensors needed to monitor T in all chambers like in RB

- Each station has 12 temperature sensors at the chambers with 60degree gap. For six station we have 72 sensors.4 humidity sensors have been installed at each disk to read the environmental humidity. Total for six stations there are 24 humidity sensors. On +Z endcap 4 temperature sensors at coolant have been installed at each station (12 sensors in total). On Z endcap, 3 sensors have been installed in this week in the accessible zones with CMS closed.Every thing has been implemented in PVSS and data can be recorded.

+ the circuitThe RH SENSOR:HIH-4000 Honeywell

The ADCCAEN A3801A 128 Channel Temp. Sensor Board(-45 C +125 C)AD592SPECIFICATIONSAD592CNACCURACY typical maximumCalibration Error @ +25C : 0.3 0.5 CError over Temperature : 0.4 1.0 CNonlinearity : 0.1 0.35C

The T sensor AD592CN Analog DevicesSensor Construction and assembly done in Pavia (circuits and plastic cases), mechanics for gas pipes done in Naples - R. Guida coordinationInstallation: Roberto, Ian, Nicolas, SashaMapping and read out: Nicolas and Giovanni

RH sensors on the disks: 4 RH sensors per disk(far, near, north, south) =24 sensorsRH sensors on the gas line (supply, return) : 4 per disk + 8 per wheel = 64 sensorsT sensors on the chambers : 12 per disk (6 ch of the inner ring + 6 ch of the outer ring) =72 sensorsT sensors on the RE-cooling pipes : +side (supply & return): 2 cooling sectors per disk =2 x2 x3 =12 sensors -side (only supply) : 1 cooling sector per disk (top) = 3 sensors (DONE NOW !)

19Endcap Pending Jobs18 chambers are single gap mode: we have to find the problematic points and solutions in P5.In one gas sector of RE+2 gas leak need to be fixed.INB labels in HV hardware in USC to be mounted.INB labels in RE-3 to be mounted.Plastic coated labels at the modules side and cables routing in cables tray.4 i2c problems need to fix.Cooling circuits to be repaired (5 chambers in RE+1, and 1 in RE-1), demanding CSC dismounting.Additional 9 T sensors on cooling water pipes in Z endcapAdditional 360 T sensors on all chambers in endcaps (now only 12 out of 72 for each disk)Endcap hardware: Spare status Hardware DescriptionQuantity in useSpare we haveMore spare we neededHigh VoltageModules (A3512N)3660MAOs(A3485)101EASY Crates(EASY3000)830Branch Controllers(A1676A)701Low VoltageModules(A3009)3623MAOs(A3486S)121(A3486)2 (A3486S)EASY crates(EASY3000S)1202128 Channel Temp. ADC Board(A3801A)4 (2 borrow from barrel)03 have been ordered

Endcap Cooling16/07/09 : A modification to the Cold Box of the magnet was made to allow a P increase (11.8 to 15.8bar) in the Primary and increase the flow by ~70%

17/07/09 : The set point was increased from 16 to 18 deg allowing the regulator valve to close so giving some control over the system

23/07/09 :Beginning of the CRAFT

27/07/09 : The set point was decreased from 18 to 17 deg, since the regulator valve remained all the time 100% open until the end of the craft

23

Temperature variations of the Secondary Cooling Circuit (week 29)Set point change from 16 -18deg CIncrease in primary pressure from 11.8 to 15.8 bar. Flow up by >70%MWGR16deg C Set point 17/07/09 ~3pm18 deg C Set point Tsup=17.9, Tret=1924

T Secondary Cooling Circuit27/07/09 ~13:0017 deg C Set point Tsup=17.2, Tret=18.4Reg Valve ~